# BCTR: Bidirectional Conditioning Transformer for Scene Graph Generation

### v2

@@ 1 Introduction
Figure 1: The motivation and pipeline paradigm of BCTR. (a) Comparison of different conditioning approaches in SGG. Unlike unidirectional conditioning, bidirectional conditioning can generate relationships through mutual feature augmentation, even when an entity is missing. However, this interaction pattern has limited generalization to unseen categories. The right panel illustrates our approach, which learns bidirectional conditioning within a semantic-aligned space, allowing it to generalize to unseen but semantically related relationships. (b) The left part shows the extraction of CLIP, entity, and predicate features from the input image. The right part demonstrates how BCTR enhances the detection of entities (e.g., ”Bag”) and predicates (e.g., ”Behind”) in the semantic-aligned feature space through bidirectional interaction.
@@
Scene Graph Generation (SGG) aims to enable computers to understand and interpret images by detecting objects and identifying the relationships between them. This process generates structured relationships in the form of triplets (object-predicate-subject). SGG has significant potential for a wide range of downstream applications, including visual question answering [25], image captioning [32], and text-to-image retrieval [11].
@@
Previous SGG studies can be categorized into two-stage and one-stage approaches. Two-stage methods separate SGG into entity and predicate detection stages, generating $\mathcal{O}(N^{2})$ relationship candidates, which demands significant computational resources. Inspired by fully convolutional one-stage object detection [1, 27], one-stage SGG methods detect relationships directly from image features [23, 30], improving detection efficiency by avoiding the consideration of all possible pairs. However, these methods lack explicit entity modeling and face challenges with complex relationships within images. To address this, recent one-stage methods [18, 26] condition predicate prediction on entity features to enhance task performance, while their fixed unidirectional dependence often yields suboptimal results, as entity detection does not benefit from predicate information.
@@
To address this limitation, we propose a bidirectional conditioning factorization in a semantic-aligned space for SGG, enabling efficient and generalizable interactions between entities and predicates, as illustrated in Fig. 1 (a). Unlike unidirectional conditioning, bidirectional conditioning can predict correct relationships by leveraging feature interactions to infer missing entity/predicate nodes (e.g., Bench). However, due to the long-tail distribution of the SGG dataset, the learned interactions in an unregularized feature space struggle to generalize to unseen but semantically related relationships (e.g., Man-Laying on-Sofa). The core insight of this paper is that learning bidirectional interaction patterns within a semantic-aligned feature space significantly enhances SGG model performance. The benefits lay in two folds. First, bidirectional interaction learning can be facilitated in the semantic feature space during training. Second, the learned interaction patterns can effectively generalize to unseen relationships during inference.
@@
  • We develop an end-to-end SGG model BCTR to implement our factorization. Specifically, BCG is designed to augment the feature spaces through bidirectional attention mechanisms, while RFA is introduced to regularize the feature space with VLPMs for facilitating BCG learning interaction patterns effectively.
@@ 2 Related Work
Conditional Dependencies in One-stage SGG Inspired by one-stage detection methods [27, 1], previous studies have designed Relation Affinity Fields [23] or used query-based detection [30] for one-stage SGG. However, these methods perform poorly as they do not utilize entity detection information [18]. Recent works [26, 4, 9] have incorporated entity features to improve relationship detection. Nonetheless, these methods either condition predicates on entities [18, 26] or condition entities on predicates [6], resulting in unidirectional dependencies that limit feature interaction and thus restrict performance gains. Khandelwal et al. [12] argue that SGG can benefit from dynamic conditioning on the image. However, their predicate remains unidirectionally conditioned on the entity, and performance improvement relies on the loss function rather than network architecture. In contrast, our proposed BCG introduces internal bidirectional dependencies, iteratively enhancing features in each round. Additionally, the RFA module enables BCG to learn feature interactions within a semantic-aligned space, improving the generalization of learned interaction patterns to semantically related relationships.
@@
Message Passing in SGG Message passing aims to boost SGG performance through interactions between predictions and their context. Zhu et al. [38] categorize it into local (within a triplet) and global (across all elements) message passing. They observe that prediction structures influence the types of message passing: triplet set-based methods [20] utilize local message passing, while chain [31], tree [28], and fully-connected graph [17] structures enable global message passing. Global message passing mitigates local ambiguities by integrating contextual information. In contrast to previous methods, we design the BiAtt-based BCG to achieve global message passing within a set structure, leveraging implicit global connections for more efficient information exchange.
Figure 2: Overview of the BCTR. (a) Visual and entity features are extracted from the input image. (b) The compositional predicate and entity queries are iteratively updated through the proposed BCG. (c) During training, the output features from various decoders are regularized by the RFA. Final predictions are generated from these distilled features.
@@
External Knowledge in SGG Enhancing scene graph generation with external knowledge is a vital research direction [38]. Previous approaches [33, 14] typically extract statistical information from textual sources like Wikipedia, but such data often fails to capture the complex patterns of commonsense, leading to limited learning improvements [21]. Gu et al. [7] propose a knowledge-based module that refines features by reasoning over a collection of commonsense knowledge retrieved from ConceptNet. However, Zareian et al. [34] noted that incomplete and inaccurate commonsense can hinder task performance. In contrast, BCTR leverages visual-language pre-trained models as a source of external knowledge. By embedding this knowledge through feature distillation, our model learns richer feature representations for scene graph generation, achieving superior performance on complex and unseen relationships.
@@ 3 Problem Formulation
Previous one-stage SGG methods typically assume a unidirectional information flow, such as $I\rightarrow s,o\rightarrow p$ [4], or $I\rightarrow p\rightarrow s,o$ [6], limiting mutual benefits between the two predictions. $I$, $o$, $s$, and $p$ are the abbreviation of the image, object, subject and predicate, respectively. Teng et al. [30] formulate SGG as $I\rightarrow p,s,o$ to facilitate feature interaction. However, optimizing over the massive compositional triplet space is challenging. This paper proposes a novel factorization for SGG, as shown in Eq. 1.
$Pr(E,P|I)=Pr(\hat{E},\widetilde{P}|I)\cdot Pr(E,P|\hat{E},\widetilde{P}).$ (1)
where $E$ and $P$ represent the entity and predicate estimates, respectively. The first term reflects $I\rightarrow s,o$ and $I\rightarrow p$, avoiding the optimization issues caused by the large compositional space. The second term enforces bidirectional dependencies $p\leftrightarrow s,o$, allowing the two predictions to benefit from each other.
After acquiring the predictions of entities and predicates, we formulate SGG as a bipartite graph construction task based on previous work [18]. Specifically, the predictions of entities and predicates from image form two node sets $\mathcal{V}_{e}$ and $\mathcal{V}_{p}$, respectively. Two directional edge sets $\mathcal{E}_{ep}$ and $\mathcal{E}_{pe}$ are used to connect these node sets, representing the connections from entities to predicates and vice versa. The bipartite graph is then represented as $\mathcal{G}_{b}=\{\mathcal{V}_{e},\mathcal{V}_{p},\mathcal{E}_{ep},\mathcal{E}_{pe}\}$, from which the scene graph of the image can be extracted.
@@ 4 Method
This section presents BCTR for implementing the factorization in Eq. 1. The overview of BCTR is shown in Fig. 2, which comprises two main modules: BCG and RFA. We detail these components in the sequel.
@@ 4.1 Feature Extraction
Inspired by the previous one-stage detection method DETR [1], we utilize a CNN and Transformer to extract features $V\in\mathbb{R}^{w\times h\times c}$ from the input image $I$, where $w$, $h$, and $c$ represent the width, height, and feature channels, respectively. Since previous one-stage SGG methods have shown that models struggle to capture predicates directly from image features without using intermediate information (e.g., entity features), we further extract entity features $V_{e}\in\mathbb{R}^{N_{e}\times c}$ from image feature $V$ as auxiliary features, as shown in the following:
$V_{e}=f_{e}(V,Q_{e}),$ (2)
where $f_{e},Q_{e}\in\mathbb{R}^{N_{e}\times c}$ represent the transformer-based decoder and the learnable queries, respectively, where $N_{e}$ denoting the number of queries.
@@ 4.2 Bidirectional Conditioning Generator
This subsection details the Bidirectional Conditioning Generator, as shown in Fig. 2. BCG comprises two interactive branches that take visual features $V$ as input and output augmented entity features and compositional predicate features, respectively. To improve performance, we introduce the iterative improvement mechanism into BCG. Specifically, we factorize the conditional distribution of entity and predicate according to Eq. 1, which are as follows:
$Pr(E^{t}|I,E^{t-1},\widetilde{P}^{t})=Pr(\hat{E}^{t}|I,E^{t-1})\cdot Pr(E^{t}|\hat{E}^{t},\widetilde{P}^{t}),$
$Pr(P^{t}|I,P^{t-1},\hat{E}^{t})=Pr(\widetilde{P}^{t}|I,P^{t-1},\hat{E}^{t})\cdot Pr(P^{t}|\hat{E}^{t},\widetilde{P}^{t}),$
@@
where $E^{t}$ and $P^{t}$ represent the entity and predicate estimates at phase $t$, respectively, while $\widetilde{}$ and $\hat{}$ denote the intermediate estimates at each phase. The first terms of the two equations predict the intermediate estimates $\widetilde{P}^{t}$ and $\hat{E}^{t}$ based on the previous estimates and the image, corresponding to the first term of Eq. 1. The second terms establish bidirectional dependencies between entities and predicates, corresponding to the second term of Eq. 1. At each phase $t$, these estimates will be updated through the corresponding decoder layer. The implementation details of Eq. 3 and Eq. 4 are introduced as follows.
@@
Inspired by previous work [18], we initialize $Q_{p}^{0}$ randomly to generate the compositional predicate queries $Q_{com}^{t-1}$ for the current scene, implemented as follows:
$Q_{com}^{t-1}=A(q=Q_{p}^{0},k=V_{e},v=V_{e}),$ (5)
@@
where $A$ denotes the attention block, and $q,k,v$ represent the query, key, and value of the attention network, respectively. $Q_{com}^{t-1}$ serves as the input to the predicate branch at phase $t$, consisting of three sub-queries: $Q_{io}^{t-1}$, $Q_{p}^{t-1}$, and $Q_{is}^{t-1}$. $Q_{e}^{t-1}$ is the input to the entity branch at phase $t$, initialized from $V_{e}$. At step $t$, $Q_{p}^{t-1}$ and $Q_{e}^{t-1}$ are updated with $V$ via a cross-attention module, defined as follows:
$\hat{Q}_{e}^{t}=A(q=Q_{e}^{t-1},k=V,v=V),$ (6)
$\hat{Q}_{p}^{t}=A(q=Q_{p}^{t-1},k=V,v=V).$ (7)
@@
The Eq. 6 corresponds to the first term of Eq. 3. Then, $Q_{io}^{t-1}$ and $Q_{is}^{t-1}$ are updated with entity features $\hat{Q}_{e}^{t}$ through cross-attention. This step aims to identify entity pairs that match the corresponding predicates from the current entity detection. The process is implemented as follows:
$Q_{io}^{t}=A(q=Q_{io}^{t-1},k=\hat{V}_{e}^{t},v=\hat{V}_{e}^{t}),$ (8)
$Q_{is}^{t}=A(q=Q_{is}^{t-1},k=\hat{V}_{e}^{t},v=\hat{V}_{e}^{t}),$ (9)
@@
where $\hat{V}_{e}^{t}=V_{e}+\lambda\,\text{Norm}(\hat{Q}_{e}^{t})$. After updating these queries with the corresponding decoder, $\hat{Q}_{p}^{t}$ is further augmented with the updated indicator queries to adjust the predicate distribution, computed as follows:
$\widetilde{Q}_{p}^{t}=(\hat{Q}_{p}^{t}+(Q_{io}^{t}+Q_{is}^{t})\cdot W_{i})\cdot W_{p}.$ (10)
@@
Eq. 7 to Eq. 10 correspond to the first term of Eq. 4, where $W_{i}$ and $W_{p}$ represent transformation matrices. The indicator queries $Q_{io}^{t}$ and $Q_{is}^{t}$ are computed from entity features, enhancing the predicate query $\widetilde{Q}_{p}^{t}$ based on current entity detections. However, entity detection does not yet leverage information from predicates. To address this, we introduce a bidirectional attention module to establish conditional dependencies between entities and predicates, enabling mutual augmentation of entity and predicate features. This module is implemented as follows:
$Q_{e}^{t}=A(q=\hat{Q_{e}^{t}},k=\widetilde{Q}_{p}^{t},v=\widetilde{Q}_{p}^{t}),$ (11)
$Q_{p}^{t}=A(q=\widetilde{Q}_{p}^{t},k=\hat{Q_{e}^{t}},v=\hat{Q_{e}^{t}}).$ (12)
@@
Eq. 11 and Eq. 12 correspond to the second terms of Eq. 3 and Eq. 4, respectively. After the bidirectional interaction, the updated queries $Q_{p}^{t}$, $Q_{io}^{t}$, $Q_{is}^{t}$ and $Q_{e}^{t}$ are used as inputs for the next phase. Through multi-stage iterative refinement, the bidirectional interaction between entities and predicates is progressively enhanced. At the end of the iterations, the final queries $Q_{com}^{\text{end}}$ and $Q_{e}^{\text{end}}$ are fed into the corresponding Multi-Layer Perceptrons (MLPs) to predict the entity and predicate distributions, respectively.

### v3

@@ 1 Introduction
Figure 1: The motivation and pipeline of BCTR. (a) Comparison of conditioning approaches in SGG. While bidirectional conditioning generates relationships through mutual feature augmentation even when an entity is missing, it struggles to generalize to unseen categories. Our approach learns bidirectional conditioning in a semantic-aligned space, achieving better generalization to unseen but semantically related relationships. (b) The left panel illustrates the extraction of CLIP, entity, and predicate features from the input image, while the right panel shows how BCTR enhances entity (e.g., "Bag") and predicate (e.g., "Behind") detection through bidirectional interaction in the semantic-aligned space.
@@
Scene Graph Generation (SGG) aims to enable computers to understand and interpret images by detecting objects and identifying the relationships between them. This process generates structured relationships in the form of triplets (object-predicate-subject). SGG has significant potential for a wide range of applications, including visual question answering [33], image captioning [41], and text-to-image retrieval [12].
@@
Previous SGG studies can be categorized into two-stage and one-stage approaches. Two-stage methods separate SGG into entity and predicate detection stages, generating $\mathcal{O}(N^{2})$ relationship candidates, which demands significant computational resources. Inspired by one-stage object detection [1, 35], one-stage SGG methods detect relationships directly from image features [28, 38], improving efficiency by avoiding the consideration of all possible pairs. However, these methods lack explicit entity modeling and face challenges with complex relationships within images. To address this, recent one-stage methods [20, 34] condition predicate prediction on entity features to enhance task performance, while their fixed unidirectional dependence often yields suboptimal results, as entity detection does not benefit from predicate information.
@@
To address this limitation, we propose a bidirectional conditioning factorization in a semantic-aligned space for SGG, enabling efficient and generalizable interactions between entities and predicates, as illustrated in Fig. 1 (a). Unlike unidirectional conditioning, bidirectional conditioning can predict correct relationships by leveraging feature interactions to infer missing entity/predicate nodes (e.g., Bench). However, due to the long-tail distribution of the SGG dataset, the learned interactions in an unregularized feature space struggle to generalize to unseen but semantically related relationships (e.g., Man-Laying on-Sofa). The core insight of this paper is that learning bidirectional interaction patterns within a semantic-aligned feature space significantly enhances SGG model performance. The benefits lie in two folds. First, bidirectional interaction learning can be facilitated in the semantic feature space during training. Second, the learned interaction patterns can effectively generalize to unseen relationships during inference.
@@
  • We present an end-to-end SGG model BCTR to implement our factorization. Specifically, BCG is designed to augment the feature spaces through bidirectional attention mechanisms.
  • We introduce RFA to regularize the feature spaces with VLPMs while preserving diversity for specific tasks. RFA facilitates BCG in learning interaction patterns from long-tail datasets and boosts model performance on tail categories.
@@
The rest of this paper is organized as follows. We briefly review related works in Section 2. In Section 3, we formulate the problem. We introduce our proposed method in Section 4. We then present a variety of experimental results and analyses in Section 5. Finally, Section 6 concludes this paper.
@@ 2 Related Work
2.1 Conditional Dependencies in One-stage SGG
The previous SGG studies can be divided into two categories: two-stage SGG and one-stage SGG. Two-stage SGG methods [22] perform entity detection first and then predict the predicate between each pair. Although these methods can capture each possible relationship in the scenario, the predictions are time-consuming and suffer from trivial outputs. In contrast, one-stage SGG methods [28, 38] directly detect predicates or relationships from image features, avoiding the combination of entities to improve detection efficiency. Inspired by one-stage detection methods [35, 1], previous SGG studies have designed Relation Affinity Fields [28] or used query-based detection [38] to directly predict triplets from images. However, these methods perform poorly as they do not utilize entity detection information [20].
@@ 2.1 Conditional Dependencies in One-stage SGG
Recent works [34, 4, 10, 9] have explicitly incorporated entity features into relationship detection to improve relationship detection. Nonetheless, these methods either condition predicates on entities [20, 34] or condition entities on predicates [6], resulting in unidirectional dependencies that limit feature interaction and thus restrict performance gains. Khandelwal et al. [14] argue that SGG can benefit from dynamic conditioning on the image. By implicitly through a novel joint loss and explicitly through cross-decoder and layer-wise attention mechanisms, their predictions are effectively refined from previous estimates. However, their predicate remains unidirectionally conditioned on the entity, and performance improvement relies on the loss function rather than network architecture. In contrast, our proposed BCG introduces internal bidirectional dependencies, iteratively enhancing features in each round. Additionally, the RFA module enables BCG to learn feature interactions within a semantic-aligned space, improving the generalization of learned interaction patterns to semantically related relationships.
2.2 Message Passing in SGG
@@
Message passing aims to boost SGG performance through interactions between predictions and their context. Zhu et al. [50] categorize it into local (within a triplet) and global (across all elements) message passing. They observe that prediction structures influence the types of message passing: triplet set-based methods [23] utilize local message passing, while chain [40], tree [37], and fully-connected graph [22] structures enable global message passing. Global message passing mitigates local ambiguities by integrating contextual information. In contrast to previous methods, we design the BiAtt-based BCG to achieve global message passing within a set structure, leveraging implicit global connections for more efficient information exchange. Moreover, previous message passing generally deploys on two-stage methods [22] [40], while the graph generation doesn’t influence the entity detections [14]. Thus, their feature interactions can’t be used to optimize entity feature extraction during training. Different from them, BCTR is a one-stage SGG method. The end-to-end mechanism enables the designed feature interaction (BCG) to facilitate the learning of entity detection.
2.3 External Knowledge in SGG
@@
Enhancing SGG performance with external knowledge is a crucial research topic [50]. Previous SGG works [42, 16] usually extract statistical information from external textual knowledge (e.g., Wikipedia, ConceptNet). However, statistical information such as co-occurrence frequency is limited in revealing the complex, structured patterns of commonsense, which may lead to poor learning improvement [25]. Generally, using external knowledge to guide feature refinement is another effective approach. Gu et al. [8] proposed a knowledge-based module that improves the feature refinement procedure by reasoning over a collection of commonsense knowledge retrieved from ConceptNet. Zhan et al. [45] introduced a novel multi-modal feature-based undetermined relationship learning network. Unfortunately, Zareian et al. [43] point out that incomplete and inaccurate external commonsense tends to limit task performance. Unlike previous works, BCTR uses VLPMs as a source of external knowledge. By embedding this knowledge through feature distillation, our model learns richer feature representations for scene graph generation, demonstrating superior performance on complex and unseen relationships.
2.4 Language Prior for Long-Tail Problem
The long-tail distribution presents a significant challenge in visual relationship recognition, leading to biased predictions in SGG methods [50]. The SGG community has developed numerous unbiased SGG methods, with one widely used approach being re-balancing. Re-balancing aims to mitigate the negative impact of the long-tail distribution through re-weighting or re-sampling [5, 22]. However, these methods are limited to open-set tasks due to their reliance on statistical priors of datasets. Chang et al. [2] argue that the long-tail problem can be addressed with language priors. Lu et al. [29] trained two models, one of which includes a language prior module that projects semantic-like relationships into a more compact embedding space, enabling the model to infer similar visual relationships (e.g., "person riding an elephant" from "person riding a horse"). Their work demonstrates the potential of solving the long-tail problem by aligning the image and language spaces. It is worth noting that language priors are scalable to open-world scenarios compared to statistical priors. However, the enhancement of using language priors to address the long-tail problem is affected by the poor performance of multi-modal feature alignment.
Recently, VLPMs have made breakthroughs in multi-modal tasks [7], achieving superior performance in aligning image and language features. By distilling features from VLPMs, downstream classifiers can recognize more objects and improve detection results on tail categories [24, 10]. However, research on using language priors in one-stage SGG is still lacking, and current methods fail to address the long-tail problem explicitly [18]. Different from previous works, our proposed RFA aligns the visual feature space of SGG with the pre-trained language space to solve the long-tail problem. To our knowledge, RFA is the first method to use a multi-modal model to address data imbalance in SGG, suggesting potential for open-set relation detection.
@@ 3 Problem Formulation
Previous one-stage SGG methods typically assume a unidirectional information flow, such as $\mathbf{I}\rightarrow\mathbf{s},\mathbf{o}\rightarrow\mathbf{p}$ [4], or $\mathbf{I}\rightarrow\mathbf{p}\rightarrow\mathbf{s},\mathbf{o}$ [6], limiting mutual benefits between the two predictions. Here, $\mathbf{I}$, $\mathbf{o}$, $\mathbf{s}$, and $\mathbf{p}$ are the abbreviations for image, object, subject and predicate, respectively. Teng et al. [38] formulate SGG as $\mathbf{I}\rightarrow\mathbf{p},\mathbf{s},\mathbf{o}$ to facilitate feature interaction. However, optimizing over the massive compositional triplet space is challenging. This paper proposes a novel factorization for SGG, as shown in Eq. 1.
$\Pr(\mathbf{E},\mathbf{P}|\mathbf{I})=\Pr(\hat{\mathbf{E}},\widetilde{\mathbf{P}}|\mathbf{I})\cdot\Pr(\mathbf{E},\mathbf{P}|\hat{\mathbf{E}},\widetilde{\mathbf{P}}).$ (1)
where $\mathbf{E}$, $\mathbf{P}$ represent the entity and predicate estimates, respectively. $\widetilde{}$ and $\hat{}$ denote the intermediate estimates. The first term reflects $\mathbf{I}\rightarrow\mathbf{s},\mathbf{o}$ and $\mathbf{I}\rightarrow\mathbf{p}$, avoiding the optimization issues caused by the large compositional space. The second term enforces bidirectional dependencies $\mathbf{p}\leftrightarrow\mathbf{s},\mathbf{o}$, allowing the two predictions to benefit from each other.
After acquiring the predictions of entities and predicates, we formulate SGG as a bipartite graph construction task based on previous work [20]. Specifically, the predictions of entities and predicates from the image form two node sets $\mathcal{V}_{e}$ and $\mathcal{V}_{p}$, respectively. Two directional edge sets $\mathcal{E}_{ep}$ and $\mathcal{E}_{pe}$ are used to connect these node sets, representing the connections from entities to predicates and vice versa. The bipartite graph is then represented as $\mathcal{G}_{b}=\{\mathcal{V}_{e},\mathcal{V}_{p},\mathcal{E}_{ep},\mathcal{E}_{pe}\}$, from which the scene graph of the image can be extracted.
@@ 4 Method
Figure 2: Overview of the BCTR. (a) Visual and entity features are extracted from the input image. (b) The entity-aware predicate queries and entity queries are iteratively updated through the proposed BCG. (c) During training, the output features from various decoders are regularized by the RFA. Final predictions are generated from these distilled features.
This section presents BCTR for implementing the factorization in Eq. 1. The overview of BCTR is shown in Fig. 2, which comprises two main modules: BCG and RFA. We detail these components in the sequel. First, we introduce the backbone for feature extraction. Next, we detail BCG, which provides mutual dependence for entity and predicate through BiAtt. RFA is then introduced to regularize the feature spaces with VLPMs and facilitate BCG learning interaction patterns from long-tail datasets. Then, we describe graph assembling, aiming to construct the scene graph from entity and predicate results. Finally, model training and inference are detailed.
@@ 4.1 Feature Extraction
Inspired by the previous one-stage detection method DETR [1], we utilize a CNN and Transformer to extract features $\mathbf{V}\in\mathbb{R}^{w\times h\times c}$ from the input image $\mathbf{I}$, where $w$, $h$, and $c$ represent the width, height, and feature channels, respectively. Since previous one-stage SGG methods have shown that models struggle to capture predicates directly from image features without using intermediate information (e.g., entity features), we further extract entity features $\mathbf{V}_{e}\in\mathbb{R}^{N_{e}\times c}$ from image feature $\mathbf{V}$ as auxiliary features, as shown in the following:
$\mathbf{V}_{e}=\mathcal{F}_{e}(\mathbf{V},\mathbf{Q}_{e}),$ (2)
where $\mathcal{F}_{e},\mathbf{Q}_{e}\in\mathbb{R}^{N_{e}\times c}$ represent the transformer-based decoder and the learnable queries, respectively, and $N_{e}$ denotes the number of queries.
@@ 4.2 Bidirectional Conditioning Generator
This subsection details the Bidirectional Conditioning Generator, as shown in Fig. 2. BCG comprises two interactive branches that take visual features $\mathbf{V}$ as input and output augmented entity features and entity-aware predicate features, respectively. To improve performance, we introduce the iterative improvement mechanism into BCG. Specifically, we factorize the conditional distribution of entity and predicate according to Eq. 1, which are as follows:
$\Pr(\mathbf{E}^{t}|\mathbf{I},\mathbf{E}^{t-1},\widetilde{\mathbf{P}}^{t})=\Pr(\hat{\mathbf{E}}^{t}|\mathbf{I},\mathbf{E}^{t-1})\cdot\Pr(\mathbf{E}^{t}|\hat{\mathbf{E}}^{t},\widetilde{\mathbf{P}}^{t}),$ (3)
$\Pr(\mathbf{P}^{t}|\mathbf{I},\mathbf{P}^{t-1},\hat{\mathbf{E}}^{t})=\Pr(\widetilde{\mathbf{P}}^{t}|\mathbf{I},\mathbf{P}^{t-1},\hat{\mathbf{E}}^{t})\cdot\Pr(\mathbf{P}^{t}|\hat{\mathbf{E}}^{t},\widetilde{\mathbf{P}}^{t}),$ (4)
@@
where $\mathbf{E}^{t}$ and $\mathbf{P}^{t}$ represent the entity and predicate estimates at phase $t$. $\widetilde{}$ and $\hat{}$ denote the intermediate estimates. The first terms of the two equations predict the intermediate estimates $\widetilde{\mathbf{P}}^{t}$ and $\hat{\mathbf{E}}^{t}$ based on the previous estimates and the image, corresponding to the first term of Eq. 1. The second terms establish bidirectional dependencies between entities and predicates, corresponding to the second term of Eq. 1. At each phase, these estimates will be updated through the corresponding decoder layer. The implementations of Eq. 3 and Eq. 4 are introduced as follows.
@@
Inspired by previous work [20], we initialize $\mathbf{Q}_{p}^{0}$ randomly to generate the entity-aware predicate queries $\mathbf{Q}_{com}^{t-1}$ for the current scene, implemented as follows:
$\mathbf{Q}_{com}^{t-1}=\mathcal{A}(q=\mathbf{Q}_{p}^{0},k=\mathbf{V}_{e},v=\mathbf{V}_{e}),$ (5)
@@
where $\mathcal{A}$ denotes the attention block, and $q,k,v$ represent the query, key, and value of attention networks, respectively. $\mathbf{Q}_{com}^{t-1}$ serves as the input to the predicate branch at phase $t$, consisting of three sub-queries: $\mathbf{Q}_{io}^{t-1}$, $\mathbf{Q}_{p}^{t-1}$, and $\mathbf{Q}_{is}^{t-1}$. $\mathbf{Q}_{e}^{t-1}$ is the input to the entity branch at phase $t$, initialized from $\mathbf{V}_{e}$. At phase $t$, $\mathbf{Q}_{p}^{t-1}$ and $\mathbf{Q}_{e}^{t-1}$ are updated with $\mathbf{V}$ via a cross-attention module, defined as follows:
$\hat{\mathbf{Q}}_{e}^{t}=\mathcal{A}(q=\mathbf{Q}_{e}^{t-1},k=\mathbf{V},v=\mathbf{V}),$ (6)
$\hat{\mathbf{Q}}_{p}^{t}=\mathcal{A}(q=\mathbf{Q}_{p}^{t-1},k=\mathbf{V},v=\mathbf{V}).$ (7)
@@
The Eq. 6 corresponds to the first term of Eq. 3. Then, $\mathbf{Q}_{io}^{t-1}$ and $\mathbf{Q}_{is}^{t-1}$ are updated with entity features $\hat{\mathbf{Q}}_{e}^{t}$ through cross-attention. This step aims to identify entity pairs that match the corresponding predicates from the current entity detection. The process is implemented as follows:
$\mathbf{Q}_{io}^{t}=\mathcal{A}(q=\mathbf{Q}_{io}^{t-1},k=\hat{\mathbf{V}}_{e}^{t},v=\hat{\mathbf{V}}_{e}^{t}),$ (8)
$\mathbf{Q}_{is}^{t}=\mathcal{A}(q=\mathbf{Q}_{is}^{t-1},k=\hat{\mathbf{V}}_{e}^{t},v=\hat{\mathbf{V}}_{e}^{t}),$ (9)
@@
where $\hat{\mathbf{V}}_{e}^{t}=\mathbf{V}_{e}+\lambda\,\text{Norm}(\hat{\mathbf{Q}}_{e}^{t})$. After updating these queries with the corresponding decoder, $\hat{\mathbf{Q}}_{p}^{t}$ is further augmented with the updated indicator queries to adjust the predicate distribution, computed as follows:
$\widetilde{\mathbf{Q}}_{p}^{t}=(\hat{\mathbf{Q}}_{p}^{t}+(\mathbf{Q}_{io}^{t}+\mathbf{Q}_{is}^{t})\cdot\mathbf{W}_{i})\cdot\mathbf{W}_{p}.$ (10)
@@
Eq. 7 to Eq. 10 correspond to the first term of Eq. 4, where $\mathbf{W}_{i}$ and $\mathbf{W}_{p}$ represent transformation matrices. The indicator queries $\mathbf{Q}_{io}^{t}$ and $\mathbf{Q}_{is}^{t}$ are computed from entity features, enhancing the predicate query $\widetilde{\mathbf{Q}}_{p}^{t}$ based on current entity detections. However, entity detection does not yet leverage information from predicates. To address this, we introduce a bidirectional attention module to establish conditional dependencies between entities and predicates, enabling mutual augmentation of entity and predicate features. This module is implemented as follows:
$\mathbf{Q}_{e}^{t}=\mathcal{A}(q=\hat{\mathbf{Q}}_{e}^{t},k=\widetilde{\mathbf{Q}}_{p}^{t},v=\widetilde{\mathbf{Q}}_{p}^{t}),$ (11)
$\mathbf{Q}_{p}^{t}=\mathcal{A}(q=\widetilde{\mathbf{Q}}_{p}^{t},k=\hat{\mathbf{Q}}_{e}^{t},v=\hat{\mathbf{Q}}_{e}^{t}).$ (12)
@@
Eq. 11 and Eq. 12 correspond to the second terms of Eq. 3 and Eq. 4, respectively. After the bidirectional interaction, the updated queries $\mathbf{Q}_{p}^{t}$, $\mathbf{Q}_{io}^{t}$, $\mathbf{Q}_{is}^{t}$, and $\mathbf{Q}_{e}^{t}$ are used as inputs for the next phase. Through multi-stage iterative refinement, the bidirectional interaction between entities and predicates is progressively enhanced. At the end of the iterations, the final queries $\mathbf{Q}_{com}^{\text{end}}$ and $\mathbf{Q}_{e}^{\text{end}}$ are fed into the corresponding Multi-Layer Perceptrons (MLPs) to predict the entity and predicate distributions, respectively.